# Entanglement in the Schwinger effect

### v1

@@ III.4 Example: two-mode squeezing for different input states
The underlying canonical transformation associated with bosonic particle creation in quantum field theory in curved spacetime, such as the Schwinger effect [Schwinger:1951nm], the Hawking [Hawking:1974] and Unruh effects [Unruh:1976], and particle creation in cosmology [Parker:1968mv, Grishchuk:1990], is a two-mode squeezing. It is, thus, essential to review this transformation under the Gaussian state formalism and apply the tools laid down in the previous subsection to quantify entanglement for different input states of interest such as the vacuum, thermal, and squeezed states. All of them are non-displaced, and hence, we will consider $\bm{\mu}=\bm{0}$ for the remaining of the paper.
@@
The example of the thermal input highlights a crucial lesson: the non-classical signal (entanglement) of genuine quantum processes can be lost due to noisy inputs, hence masking the “quantumness” of the process. Although one could overcome this challenge by increasing the two-mode squeezing parameter $r$, in several setups, there are physical constraints limiting the value of $r$. For example, in the Hawking effect, $r$ depends on the Hawking temperature which is very low for astrophysical black holes. In the Schwinger effect, $r$ depends on the strength of the electric field, which is also physically constrained (see discussion in Section IV.3). Hence, to revive the quantum output of the process, one may turn into modifying the input state. Introducing non-classicalities, i.e. features of states with strong departure from their classical field theory description, in the input can amplify the amount of entanglement produced during the two-mode squeezing process. One such non-classical state is the single-mode squeezed state. [^12] Not to be confused with a two-mode squeezed state. These are quantum states with uncertainty in one of their quadrature operators, i.e. position or momentum, below the uncertainty of the vacuum state, hence “squeezed”, whereas the conjugate quadrature exhibits higher-than the vacuum uncertainty, hence “antisqueezed”. Since “anticommutativity” is a characteristic of quantum mechanics, giving rise to the uncertainty principle without a classical physics analogue, the antisqueezing of a quadrature pair which amplifies the uncertainty in a pair of quadratures, makes this states colloquially referred to as “non-classical”. It is also worth mentioning that single-mode squeezed states are not entangled; the squeezing is an operation on a single mode (degree of freedom). The idea of introducing non-classical inputs to enhance the amount of entanglement in particle creation processes was introduced in [Agullo:2021vwj, Brady:2022ffk] where it was shown that the use of a single-mode squeezed state in the input can amplify the amount of entanglement produced by the two-mode squeezing process. In this paper we apply the same protocol for the Schwinger effect. The reason for using a single-mode squeezed state in the input instead of, for example other non-classical inputs, such as a two-mode squeezed state is to avoid artificially introducing entanglement in the system.
@@ IV.1.1 Number of created particles
where $\bm{S}$ is the two-mode squeezer matrix given in eq. (56) with $r$ and $\varphi$ fixed by the Bogoliubov coefficients we computed in Section II $(\alpha_{I}\equiv\alpha_{\lambda_{I}},\beta_{I}\equiv\beta_{\lambda_{I}})$. More explicitly [^13] The parametrization of the two-mode squeezer assumes $\alpha_{I}$ is real, while the $\alpha_{I}$ we computed for the Schwinger effect is not. This is not a problem, since we can always perform a rotation that changes the phase of the Bogoliubov coefficients. This transformation does not affect particle number or entanglement [Birrell:1982ix].
@@
Note that $|\beta_{I}|^{2}$ is independent of $k_{z}$, which is because we are working with a constant electric field, so modes with any value of $k_{z}$ will eventually be produced. This means that the energy spectrum behaves, at high $k_{z}$, as $n(\omega)\sim\omega^{2}$ (without exponential suppression in $k_{z}$), which gives a divergent result for the average energy density, since we are considering particles that are created for an infinitely long time. [^14] For a time-dependent electric field we would obtain a $k_{z}$-dependent $\beta$ coefficient that vanishes at $|k_{z}|\to+\infty$. This is, for instance, the case for a pulsed electric field (see for example Ref. [Li_2017] or Ref. [Beltran-Palau:2019bvz] in 1+1). In the literature, for the case with constant $E$, a cutoff off in the longitudinal momentum $k_{z}$, namely $0<k_{z}<eE\Delta t$ is used. This cutoff has been verified numerically in Ref. [Domcke:2019qmm]. To get rid of this problem, we can consider, for example, only particles with no momentum along the $z$ direction $k_{z}=0$ and find (in the $B=0$ case)
@@ IV.1.3 Quantum signal of the Schwinger effect
Figure 3: (a) Phase diagram-like representation of the logarithmic negativity of bosons. We clearly see that there exists a temperature $T_{c}/\omega$ above which there can’t be any entanglement because $|\beta|$ can’t be higher than 11. (b) Same for different values of squeezing, the shaded areas correspond to regions where $L_{N}>0$.
@@ IV.2.1 Number of created particles
$\rho_{I}^{\text{(out)}}=\text{Tr}_{\bar{I}}[\hat{\rho}^{(\text{out})}_{I\bar{I}}]=\begin{pmatrix}(1-f_{I})^{2}|\alpha_{I}|^{2}+f_{I}^{2}|\beta_{I}|^{2}+f_{I}(1-f_{I})&0\\ 0&(1-f_{I})^{2}|\beta_{I}|^{2}+f_{I}^{2}|\alpha_{I}|^{2}+f_{I}(1-f_{I})\end{pmatrix}\,$, (103)
and $\rho_{\bar{I}}^{\text{(out)}}=\rho_{I}^{\text{(out)}}$. The average number of particles (and of antiparticles) of a given type ($R$ or $L$) and in a given mode $I$ ($\bar{I}$) is given by
@@ IV.2.3 Quantum signal of the Schwinger effect
We proceed with logarithmic negativity to evaluate the amount of entanglement produced. For fermions, we use directly the defining equation given in (34), since the density matrix is finite, namely $L_{N}=\log_{2}(\sum_{i}|\tilde{\lambda}_{i}|)$ where $\tilde{\lambda}_{i}$ are the eigenvalues of the partially transposed density matrix $\hat{\tilde{\rho}}$. We focus directly on the out-basis. In our case, $\rho_{I\bar{I}}^{(\text{out})}$ is given in (102), and $\hat{\tilde{\rho}}_{I\bar{I}}$ is obtained as follows: if the components of the density matrix in the Fock basis $\{|ij\rangle\langle kl|\}_{i,j,k,l=0,1}$ are $\rho_{ijkl}$, then the components of $\hat{\tilde{\rho}}_{I\bar{I}}$ are $\rho_{ilkj}$. That is, we exchange the upper-right and lower-left $2\times 2$ blocks in (102) and get
$\hat{\tilde{\rho}}^{(\text{out})}=\begin{pmatrix}(1-f_{I})^{2}|\alpha_{I}|^{2}+f_{I}^{2}|\beta_{I}|^{2}&0&0&0\\ 0&f_{I}(1-f_{I})&(1-2f_{I})\,\alpha_{I}\beta_{I}&0\\ 0&(1-2f_{I})\,\alpha_{I}^{*}\beta_{I}^{*}&f_{I}(1-f_{I})&0\\ 0&0&0&(1-f_{I})^{2}|\beta_{I}|^{2}+f_{I}^{2}|\alpha_{I}|^{2}\end{pmatrix}\,$. (112)
Its eigenvalues are
$\tilde{\lambda}_{1}=(1-f_{I})^{2}|\alpha_{I}|^{2}+f_{I}^{2}|\beta_{I}|^{2}\,,\qquad\qquad\tilde{\lambda}_{3}=f_{I}(1-f_{I})+(1-2f_{I})|\alpha_{I}||\beta_{I}|\,$, (113)
$\tilde{\lambda}_{2}=(1-f_{I})^{2}|\beta_{I}|^{2}+f_{I}^{2}|\alpha_{I}|^{2}\,,\qquad\qquad\tilde{\lambda}_{4}=f_{I}(1-f_{I})-(1-2f_{I})|\alpha_{I}||\beta_{I}|\,.$
We note that $\tilde{\lambda}_{1,2,3}$ are always positive while $\tilde{\lambda}_{4}$ can be positive or negative. As we will see, only if $\tilde{\lambda}_{4}$ is negative can the pair (II~)($\tilde{I}$) be entangled. It can indeed be shown that
$\sum_{i}|\tilde{\lambda}_{i}|=\begin{cases}1,&\tilde{\lambda}_{4}\geq 0\\ 1-2\tilde{\lambda}_{4},&\tilde{\lambda}_{4}<0\end{cases}$ (114)
Therefore, the logarithmic negativity can be written as $L_{N}=\frac{1}{\ln(2)}\text{max}(0,\ln(1-2\tilde{\lambda}_{4}))$. More explicitly, in terms of $\omega_{I}$, $T$, and $|\beta_{I}|^{2}$
$L_{N}=\frac{1}{\ln(2)}\max\left(0,\ln\left(\frac{1+e^{2\omega_{I}/T}+2(e^{2\omega_{I}/T}-1)|\beta_{I}|\sqrt{1-|\beta_{I}|^{2}}}{(1+e^{\omega_{I}/T})^{2}}\right)\,.\right)$ (115)
@@
Figure 7: Phase-like diagram representation of the logarithmic negativity of fermions. We clearly see that there exists a temperature $T_{c}/\omega$ above which there can’t be any entanglement.
There is entanglement (in the sense $L_{N}>0$) if and only if
[…]
(if $\tilde{\lambda}_{4}<0$, or, equivalently, $1-2\tilde{\lambda}_{4}>1$). In terms of the temperature, we say that a pair $(I\bar{I}\,)$ is entangled if and only if
@@
$\frac{T}{\omega}<\frac{T_{c}}{\omega}=\frac{1}{\text{arcsinh}\left((1-(2|\beta_{I}|^{2}-1)^{2})^{-1/2}\right)}$ (117)
In this case, the maximum value of the log-negativity (115) is achieved at the value of $|\beta_{I}|$ that maximizes $|\beta|\sqrt{1-|\beta_{I}|^{2}}$, that is, at $|\beta_{I}|=1/\sqrt{2}$. We thus see that the system can be entangled only if the temperature is lower than a threshold value
$T_{\mathrm{max}}=\frac{\omega}{\text{arcsinh}(1)}\approx 1.13\,\omega\,.$ (118)
In terms of the electric field, the critical temperature $T_{c}$ is not a monotonously increasing function of $E$, since it peaks for $|\beta_{I}|^{2}=1/2$ and not for $|\beta_{I}|^{2}\to 1$ (that corresponds to the large $E$ limit). As we will see in the next subsection, this has very interesting consequences.
IV.2.4 Critical field for entanglement
As in the scalar case, for a fixed temperature $T<T_{\text{max}}$ we can try to find a critical value of the electric field for entanglement. In this case, however, the entanglement condition (116) is a bit more subtle, since $|\beta_{I}|$ is between two limiting values. This implies that to have entanglement between a pair of modes, the electric field must be inside the window
$E_{\text{min}}<E<E_{\text{max}}\,$, (119)
with
@@ IV.2.4 Critical field for entanglement
$E_{\text{min}}=\frac{\pi\left(m^{2}+k_{\perp}^{2}\right)}{e\ln\left(\frac{2}{1-A(\omega/T)}\right)}\,,\qquad E_{\text{max}}=\frac{\pi\left(m^{2}+k_{\perp}^{2}\right)}{e\ln\left(\frac{2}{1+A(\omega/T)}\right)}\,$, (120)
where $A(x)=\sqrt{1-1/\sinh^{2}(x)}$. Let us analyze this result. First, we note that at zero temperature $T\rightarrow 0$, $E_{\min}\rightarrow 0$, and $E_{\max}\rightarrow\infty$. That is, any value of the external field produces entanglement for any mode. Furthermore, the field that maximizes the entanglement is
$E_{*}=\frac{\pi\left(m^{2}+k_{\perp}^{2}\right)}{e\ln 2}\,.$ (121)
This critical value is independent of $T$, but gives $L_{N}>0$ if and only if $T<T_{\text{max}}$. At the threshold temperature $T=T_{\max}=\omega_{I}/\operatorname{arcsinh}(1)$, the window collapses to the single field $E_{*}$. At this point $L_{N}=0$. For $T>T_{\max}$ there is no field giving entanglement. Finally, for fixed $k_{\perp}$ and $T$, there is always a value of $|k_{z}|$ above which there is entanglement. Furthermore, the bigger is $|k_{z}|$, the bigger is the distance between $E_{\text{min}}$ and $E_{\text{max}}$. In Figure 8 we represent all relevant critical fields as a function of the temperature.
Figure 8: Critical electric fields as expressed in eqs. (110) and (120) in the low-momentum limit $k_{\perp}=k_{z}=0$ (in the $\vec{B}=\vec{0}$ case). The shaded area corresponds to the region where the system is entangled.
IV.2.5 Axial anomaly

### v2

@@ III.4 Example: two-mode squeezing for different input states
The underlying canonical transformation associated with bosonic particle creation in quantum field theory in curved spacetime, such as the Schwinger effect [68], the Hawking [40] and Unruh effects [77], and particle creation in cosmology [55, 36], is a two-mode squeezing. It is, thus, essential to review this transformation under the Gaussian state formalism and apply the tools laid down in the previous subsection to quantify entanglement for different input states of interest such as the vacuum, thermal, and squeezed states. All of them are non-displaced, and hence, we will consider $\bm{\mu}=\bm{0}$ for the remaining of the paper.
@@
The example of the thermal input highlights a crucial lesson: the non-classical signal (entanglement) of genuine quantum processes can be lost due to noisy inputs, hence masking the “quantumness” of the process. Although one could overcome this challenge by increasing the two-mode squeezing parameter $r$, in several setups, there are physical constraints limiting the value of $r$. For example, in the Hawking effect, $r$ depends on the Hawking temperature which is very low for astrophysical black holes. In the Schwinger effect, $r$ depends on the strength of the electric field, which is also physically constrained (see discussion in Section IV.3). Hence, to revive the quantum output of the process, one may turn into modifying the input state. Introducing non-classicalities, i.e. features of states with strong departure from their classical field theory description, in the input can amplify the amount of entanglement produced during the two-mode squeezing process. One such non-classical state is the single-mode squeezed state. [^12] Not to be confused with a two-mode squeezed state. These are quantum states with uncertainty in one of their quadrature operators, i.e. position or momentum, below the uncertainty of the vacuum state, hence “squeezed”, whereas the conjugate quadrature exhibits higher-than the vacuum uncertainty, hence “antisqueezed”. Since “anticommutativity” is a characteristic of quantum mechanics, giving rise to the uncertainty principle without a classical physics analogue, the antisqueezing of a quadrature pair which amplifies the uncertainty in a pair of quadratures, makes this states colloquially referred to as “non-classical”. It is also worth mentioning that single-mode squeezed states are not entangled; the squeezing is an operation on a single mode (degree of freedom). The idea of introducing non-classical inputs to enhance the amount of entanglement in particle creation processes was introduced in [6, 17] where it was shown that the use of a single-mode squeezed state in the input can amplify the amount of entanglement produced by the two-mode squeezing process. In this paper we apply the same protocol for the Schwinger effect. The reason for using a single-mode squeezed state in the input instead of, for example other non-classical inputs, such as a two-mode squeezed state is to avoid artificially introducing entanglement in the system.
@@ IV.1.1 Number of created particles
where $\bm{S}$ is the two-mode squeezer matrix given in eq. (57) with $r$ and $\varphi$ fixed by the Bogoliubov coefficients we computed in Section II $(\alpha_{I}\equiv\alpha_{\lambda_{I}},\beta_{I}\equiv\beta_{\lambda_{I}})$. More explicitly [^13] The parametrization of the two-mode squeezer assumes $\alpha_{I}$ is real, while the $\alpha_{I}$ we computed for the Schwinger effect is not. This is not a problem, since we can always perform a rotation that changes the phase of the Bogoliubov coefficients. This transformation does not affect particle number or entanglement [16].
@@
Note that $|\beta_{I}|^{2}$ is independent of $k_{z}$, which is because we are working with a constant electric field, so modes with any value of $k_{z}$ will eventually be produced. This means that the energy spectrum behaves, at high $k_{z}$, as $n(\omega)\sim\omega^{2}$ (without exponential suppression in $k_{z}$), which gives a divergent result for the average energy density, since we are considering particles that are created for an infinitely long time. [^14] For a time-dependent electric field we would obtain a $k_{z}$-dependent $\beta$ coefficient that vanishes at $|k_{z}|\to+\infty$. This is, for instance, the case for a pulsed electric field (see for example Ref. [49] or Ref. [11] in 1+1). In the literature, for the case with constant $E$, a cutoff off in the longitudinal momentum $k_{z}$, namely $0<k_{z}<eE\Delta t$ is used. This cutoff has been verified numerically in Ref. [21]. To get rid of this problem, we can consider, for example, only particles with no momentum along the $z$ direction $k_{z}=0$ and find (in the $B=0$ case)
@@ IV.1.3 Quantum signal of the Schwinger effect
Figure 3: (a) Phase diagram-like representation of the logarithmic negativity of bosons. We clearly see that there exists a temperature $T_{c}/\omega$ above which there can’t be any entanglement because $|\beta|$ can’t be higher than $1$. (b) Same for different values of squeezing, the shaded areas correspond to regions where $L_{N}>0$.
@@ IV.2.1 Number of created particles
$\hat{\rho}_{I}^{\text{(out)}}=\text{Tr}_{\bar{I}}[\hat{\rho}^{(\text{out})}_{I\bar{I}}]=\begin{pmatrix}(1-f_{I})^{2}|\alpha_{I}|^{2}+f_{I}^{2}|\beta_{I}|^{2}+f_{I}(1-f_{I})&0\\ 0&(1-f_{I})^{2}|\beta_{I}|^{2}+f_{I}^{2}|\alpha_{I}|^{2}+f_{I}(1-f_{I})\end{pmatrix}\,$, (104)
and $\hat{\rho}_{\bar{I}}^{\text{(out)}}=\hat{\rho}_{I}^{\text{(out)}}$. The average number of particles (and of antiparticles) of a given type ($R$ or $L$) and in a given mode $I$ ($\bar{I}$) is given by
@@ IV.2.3 Quantum signal of the Schwinger effect
We proceed with logarithmic negativity to evaluate the amount of entanglement produced. For fermions, we use directly the defining equation given in (34), since the density matrix is finite, namely $L_{N}=\log_{2}(\sum_{i}|\tilde{\lambda}_{i}|)$ where ${|}\tilde{\lambda}_{i}{|}$ are the eigenvalues of the absolute value of the partially “time reversed” density matrix $\sqrt{\hat{\tilde{\rho}}\hat{\tilde{\rho}}^{\dagger}}$. We focus directly on the out-basis. In our case, $\rho_{I\bar{I}}^{(\text{out})}$ is given in (103), and $\hat{\tilde{\rho}}_{I\bar{I}}$ is obtained as follows: if the components of the density matrix in the Fock basis $\{|ij\rangle\langle kl|\}_{i,j,k,l=0,1}$ are $\rho_{ijkl}$, then the components of $\hat{\tilde{\rho}}_{I\bar{I}}$ are $(-1)^{\phi_{ijkl}}\rho_{ilkj}$ where $\phi_{ijkl}$ is a phase that depends on the occupation numbers $i,j,k,l$ as follows:
$\phi_{ijkl}=\frac{(i+k)\penalty 10000\ [\mathrm{mod}\penalty 10000\ 2]}{2}+(i+k)(j+l)\,.$ (113)
That is, we exchange the upper-right and lower-left $2\times 2$ blocks in (103) and give them an additional $\pm i$ factor and get
$\hat{\tilde{\rho}}^{(\text{out})}=\begin{pmatrix}(1-f_{I})^{2}|\alpha_{I}|^{2}+f_{I}^{2}|\beta_{I}|^{2}&0&0&0\\ 0&f_{I}(1-f_{I})&-i(1-2f_{I})\,\alpha_{I}\beta_{I}&0\\ 0&-i(1-2f_{I})\,\alpha_{I}^{*}\beta_{I}^{*}&f_{I}(1-f_{I})&0\\ 0&0&0&(1-f_{I})^{2}|\beta_{I}|^{2}+f_{I}^{2}|\alpha_{I}|^{2}\end{pmatrix}\,$. (114)
Its absolute value is
$|\hat{\tilde{\rho}}^{(\text{out})}|\equiv\sqrt{\hat{\tilde{\rho}}^{(\text{out})}\hat{\tilde{\rho}}^{(\text{out})\dagger}}=\begin{pmatrix}(1-f_{I})^{2}|\alpha_{I}|^{2}+f_{I}^{2}|\beta_{I}|^{2}&0&0&0\\ 0&s_{I}&0&0\\ 0&0&s_{I}&0\\ 0&0&0&(1-f_{I})^{2}|\beta_{I}|^{2}+f_{I}^{2}|\alpha_{I}|^{2}\end{pmatrix}\,$. (115)
with $s_{I}=\sqrt{f_{I}^{2}(f_{I}-1)^{2}+|\alpha_{I}|^{2}|\beta_{I}|^{2}(1-2f_{I})^{2}}$.
The eigenvalues of $|\hat{\tilde{\rho}}^{(\text{out})}|$ can then be summed to yield the (fermionic) logarithmic negativity:
$L_{N}=\log_{2}(\text{Tr}|\hat{\tilde{\rho}}^{(\text{out})}|)=\frac{1}{\ln(2)}\ln\left(f^{2}_{I}+(1-f_{I})^{2}+2\sqrt{f_{I}^{2}(1-f_{I})^{2}+(1-2f_{I})^{2}|\alpha_{I}|^{2}|\beta_{I}|^{2}}\right)\,.$ (116)
More explicitly, in terms of $\omega_{I}$, $T$ and $|\beta_{I}|^{2}$,
$L_{N}=\frac{1}{\ln(2)}\ln\left(1+e^{2\omega_{I}/T}+2\sqrt{e^{2\omega_{I}/T}+(e^{2\omega_{I}/T}-1)^{2}|\beta_{I}|^{2}(1-|\beta_{I}|^{2})}\right)-\frac{2}{\ln(2)}\ln(1+e^{\omega_{I}/T})\,.$ (117)
@@
Contrary to the bosonic case, the logarithmic negativity for fermions does not display a temperature cut-off. As discussed in Section III, this behaviour relies on using the fermionic definition of logarithmic negativity based on partial time-reversal, which provides a well-behaved entanglement quantifier for fermionic systems [71]. By contrast, if one applies the usual bosonic partial transpose to the same fermionic density matrix, one generally obtains a different quantity that can display, for example, spurious critical temperatures, reflecting the fact that the bosonic prescription is not a good entanglement measure for fermionic systems. In what follows, we will study the general behaviour of $L_{N}$ and some limiting cases.
First, we note that, at fixed $|\beta_{I}|$, the logarithmic negativity decreases as $T$ increases. In particular, for high temperatures we find that the logarithmic negativity decreases quadratically with the temperature, namely
$L_{N}=\frac{(|\beta_{I}|^{2}-1)|\beta_{I}|^{2}}{\ln(2)}\frac{\omega^{2}}{T^{2}}+\mathcal{O}\left(\frac{\omega^{4}}{T^{4}}\right)\,.$ (118)
On the other hand, the logarithmic negativity at fixed $|\beta_{I}|$ reaches its maximum at $T=0$ (or $f_{I}=0$)
$L_{N}^{(0)}=\frac{\ln(1+2|\beta_{I}|\sqrt{1-|\beta_{I}|^{2}})}{\ln(2)}\,.$ (119)
At fixed $T$ (or fixed $f_{I}$), we find that the logarithmic negativity has a maximum at $|\beta_{I}|^{2}=1/2$,
$L_{N}^{\text{(max)}}=\frac{\ln\left(2+4f_{I}(f_{I}-1)\right)}{\ln(2)}=\frac{1}{\ln(2)}\ln\left(1+\tanh^{2}\left(\frac{\omega_{I}}{2T}\right)\right)\,.$ (120)
Note that, $L_{N}^{(\text{max})}$ reaches its maximum value at $T=0$, namely $L_{N}^{\text{(max;0)}}=1$. This is expected, since entanglement is bounded by the finite local Hilbert-space dimension [70].
In addition, we also note that $L_{N}$ goes to zero for $|\beta_{I}|^{2}\to 0$ and for $|\beta_{I}|^{2}\to 1$.161616note that $L_{N}$ is invariant under the change $|\beta_{I}|^{2}\to(1-|\beta_{I}|^{2})$ When $|\beta_{I}|$ approaches either of these two limiting values, the logarithmic negativity goes as
$L_{N}=\frac{(1-2f_{I})^{2}}{f_{I}(1-f_{I})\ln(2)}\,u_{I}^{2}+\mathcal{O}(u_{I}^{4})=\frac{2}{\ln(2)}\left(\cosh\left(\frac{\omega_{I}}{T}\right)-1\right)u_{I}^{2}+\mathcal{O}(u_{I}^{4})$ (121)
where $u_{I}^{2}=|\beta_{I}|^{2}$ for the limit $|\beta_{I}|^{2}\to 0$ and $u_{I}^{2}=(1-|\beta_{I}|^{2})$ for $|\beta_{I}|^{2}\to 1$.
@@
Figure 7: Phase-like diagram representation of the logarithmic negativity of fermions. Contrarily to the bosonic case, as long as $|\beta|>0$ there exists no temperature above which the logarithmic negativity vanishes.
IV.2.4 Optimal field for entanglement
As in the scalar case, it is interesting to analyze the logarithmic negativity in terms of the electric field. In this context, we easily see that, at fixed $T$, there is a value of the electric field for which the entanglement is maximal, i.e., $L_{N}=L_{N}^{(\text{max})}$. This value is obtained for $|\beta_{I}|^{2}=1/2$. Therefore, the field that maximizes entanglement reads
$eE_{*}=\frac{\pi(m^{2}+k_{\perp}^{2})}{\ln 2}\,.$ (122)
This critical value is independent of $T$, however, it gives smaller values of $L_{N}^{(\text{max})}$ as $T$ increases. For $E>E_{*}$ or $E<E_{*}$ the fermionic logarithmic negativity is always smaller than $L_{N}^{(\text{max})}$.
IV.2.5 Finite-resolution criterion
We now explore benchmark values of the electric field and temperature leading to a value for the logarithmic negativity above a detectability threshold $\epsilon$, i.e. $L_{N}>\epsilon$. This lower bound is operationally motivated by experimental uncertainties in the construction of the density matrix. Our objective is to formulate analytical expressions that an experimenter could directly use to assess entanglement detectability, given the minimum value of logarithmic negativity that can be resolved, and determine the range of temperatures and electric fields needed for such a detection. For illustration purposes, in Appendix (C), we demonstrate how a lower bound $\epsilon$ in logarithmic negativity detection can emerge given the error $\delta$ in determining the density matrix elements.
In what follows, we will use the following condition for the logarithmic negativity
$L_{N}>\epsilon\,$, (123)
with $0<\epsilon\ll 1$ to define effective critical scales for $T$ and $E$ by solving $L_{N}=\epsilon$.
  • Fixed $E$. At fixed $|\beta_{I}|^{2}$ (or fixed $E$), the condition (123) defines an effective critical temperature,
@@ IV.2.5 Finite-resolution criterion
    $T_{c}(\epsilon)=\frac{\omega}{\ln\left(\frac{1+2^{\epsilon/2}g(\epsilon)}{1-2^{\epsilon/2}g(\epsilon)}\right)}\,,\qquad\text{with}\qquad g(\epsilon)=\sqrt{\frac{2^{\epsilon}-1}{2^{\epsilon}-1+4(1-|\beta_{I}|^{2})|\beta_{I}|^{2}}}\,.$ (124)
    Imposing the square root to be real requires $\epsilon<L_{N}^{(0)}$ (i.e., $\epsilon$ should be smaller than the maximum value allowed at fixed $|\beta_{I}|$, which is obtained in the zero temperature limit), with $L_{N}^{(0)}$ given in eq. (119). For very small $\epsilon$, we find
    $T_{c}(\epsilon)=\frac{\omega_{I}}{2}\frac{2^{L_{N}^{(0)}}-1}{\sqrt{\ln(2)\,\epsilon}}+\mathcal{O}\!\left(\epsilon\right)\,.$ (125)
    As expected, for $\epsilon\to 0$, $T_{c}\to\infty$.
  • Fixed $T$. At fixed $T$, the condition given in eq. (123), results in the following condition for the electric field strength
    $E_{-}(\epsilon)<E<E_{+}(\epsilon)\,$, (126)
    where
@@
    $eE_{\pm}(\epsilon)=\frac{\pi(m^{2}+k_{\perp}^{2})}{-\ln|\beta_{\pm}|^{2}}\,,\qquad\text{with}\qquad|\beta_{\pm}(\epsilon)|^{2}=\frac{1}{2}\pm\frac{2^{\epsilon/2}}{2}\sqrt{1-(2^{\epsilon}-1)\,\text{coth}^{2}\left(\frac{\omega}{2T}\right)}\,.$ (127)
    Imposing the square root to be real requires $\epsilon<L_{N}^{(\text{max})}$, which is equivalent to demanding that the error $\epsilon$ lies below the maximum negativity at fixed $T$, which is given in eq. (120). For $\epsilon\to 0$, one recovers $eE_{-}\to 0$ and $eE_{+}\to\infty$, consistently with $L_{N}\to 0$ as $|\beta_{I}|^{2}\to 0$ or $1$.
In this sense $T_{c}(\epsilon)$ is the maximum temperature for which the entanglement of a given mode remains detectable (i.e. $L_{N}>\epsilon$) at fixed field strength, while $E_{\pm}(T,\epsilon)$ delimit the field window where entanglement remains detectable at fixed temperature. Operationally, a claim of entanglement detection requires working below $T_{c}$ and within the window $E_{-}(\epsilon)<E<E_{+}(\epsilon)$; outside this region the state may still be weakly entangled, but below the experimental sensitivity encoded in $\epsilon$.
IV.2.6 Axial anomaly